f(x+y)=f(x)+f(y)...............∑ f(x)=84 (upto n terms......starting from x=1)...........f(1)=3.............then find n....PLS EXPLAIN THIS QUESTION >>>>.........

Grade:9

1 Answers

jitender lakhanpal
12 years ago

Dear Shivam,

 

f(x+y)=f(x)+f(y)  put x = y = 1      f(1) = 3

f(2) = 6

 put   x= 1, y = 2

f(3) = 9

similarly f(4) = 12

so this is an A.P starting from 3  and the common difference is also 3 

the 

∑ f(x)=84   starting from 1 is the sum of first n terms 

so by the formula 

84 = n/2[2a + (n-1)d]

so putting the values we get the quadratic in n

n2 + n - 56 = 0

so solving we get the value of n = 7

 

 

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

jitender